We keep the switch at position x for a sufficient time and thereafter (at t0) connect it to y, in the cercuit shown below : what amount will be measured through the resistor after 5.0 micro-seconds from t0, if v=9v, R= 10ohms and c=1 micro F
1
Expert's answer
2021-08-17T10:29:12-0400

Given:

"R=10\:\Omega"

"C=1*10^{-6}\:\rm F"

"V=9\:\rm V"

"t=5*10^{-6}\:\rm s"


The time constant

"\tau=RC=10*1*10^{-6}=10^{-5}\:\rm 1/s"

The current through the resistor

"I=\frac{V}{R}e^{-t/\tau}"

"I=\frac{9}{10}e^{-5*10^{-6}/10^{-5}}=0.55\:\rm A"

The voltage

"V_R=I*R=0.55*10=5.5\:\rm V"